jim is now three times as old as kathy.
Four years ago, jim was seven times as old as kathy then.
find their present ages.
PLEASE HELP ME SOLVE THIS!
and can you show work!

Answer by ankor@dixie-net.com(22740)

jim is now three times as old as kathy.
in equation form: j = 3k
:
Four years ago, jim was seven times as old as kathy then.
[J - 4] = 7[k - 4]
Substitute 3k for J, solve for k
3k - 4 = 7[k - 4]
3k - 4 = 7k - 28
-4 + 28 = 7k - 3k
24 = 4k
k = 24/4
k = 6 years old
:
J = 3(6) = 18 years old
:
Check: 4 years ago, J = 14 and K was 2 years old; J was 7 times K's age